# ChatQA 2: Bridging the Gap to Proprietary LLMs in Long Context and RAG Capabilities

### v1

@@ 6 Evaluation Benchmarks
We compare our model against SOTA long context models: 1) GPT-4-Turbo-2024-04-09 (128K context window) (OpenAI, 2023), 2) Qwen2-72B-Instruct (128K context window) (Alibaba-QWen, 2024), and (3) Llama-3-70B-Instruct-Gradient-262k (GradientAI, 2024).
@@
To give a comprehensive study of different context lengths, our evaluation benchmarks covers three categories, 1) long context benchmarks beyond 100K tokens, 2) medium-long context benchmarks within 32K tokens, and 3) short context benchmarks within 4K tokens. We also apply RAG when it is applicable to the downstream tasks.
6.1 Long Context Benchmarks Beyond 100K Tokens
@@
InfiniteBench (Zhang et al., 2024b) is proposed to evaluate the long context capability of LLMs over 100K sequence length. As we focus on real-world english tasks, we only take the four related tasks from the InfiniteBench, i.e. longbook summarization (En.Sum), longbook qa (En.QA), longbook multiple choice (En.MC), and longbook dialogue (En.Dia). En.Sum is a task that requires models to generate a concise summary of the given novel and is evaluated using the ROUGE-L-Sum metric (Lin, 2004). En.QA is annotated by a pipeline that ensures the questions’ necessitating of long-range dependencies and reasoning, beyond simple short passage retrieval. Aggregation reasoning and filtering reasoning are the two primary reasoning categories. F1 score is used to evaluate the quality of the answer. En.MC is annotated with the same pipeline of En.QA except that four answer choices are provided and exact matching scores are reported. En.Dia leverages movie and drama scripts from a designated online database  [^3] https://imsdb.com with long, multi-role dialogues. In this task, random instances of character names within a script are masked and the objective is to correctly identify these masked names. Exact matching score is used again to evaluate the prediction accuracy.
6.2 Medium-Long Context Benchmarks within 32K Tokens
@@
We use the long context datasets (except NarrativeQA as it is included in our training) from Xu et al. (2024) as our benchmark for medium-long datasets within 32K. There are six datasets in total, where QMSum (QM), Qasper (QASP), QuALITY (QLTY) are token from SCROLLS and HotpotQA (HQA) MuSiQue (MSQ), MultiFieldQA-en (MFQA) are token from LongBench. Following the official metrics, we report the geometric mean of ROUGE scores (i.e., ROUGE1/2/L) (Lin, 2004) for QM, the exact matching (EM) score for QLTY, and F1 scores for the remaining four datasets QASP, MSQ, HQA and MFQA.
@@ 6.3 Short Context within 4K Tokens
We use ChatRAG Bench (Liu et al., 2024) as our benchmark for short context within 4k. ChatRAG bench consists of 10 datasets and we exclude HDial as it is vasincluded in our training. Following the setup of Liu et al. (2024), for Doc2Dial (D2D), QuAC, and QReCC task with long documents, each document is divided into segments of roughly 300 words. The top 5 relevant chunks are then retrieved as context for each user question. For TopiOCQA and INSCIT, top-20 chunks were retrieved to obtain similar context length to the first three datasets. The other four datasets are CoQA, DoQA, ConvFinQA (CFQA), and SQA, which cover a wide range of domains like finance, children’s stories, literature, mid/high school exams, news, Wikipedia and etc. We use F1 score as the metric to evaluate the generations and report the average score without HDial as it is a fair zero-shot comparisons over different models.
@@ 7 Results
Figure 1: Needle in A Haystack test for $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ up to 128K context window.
@@ 7.1 Needle In A Haystack
We evaluate our $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ model on the Needle In A Haystack test (Kamradt, 2023). This synthetic task is popular for testing the long-context capability of LLMs, and can be considered as a threshold level evaluation. Figure 1 demonstrates the performance of our model with up to 128K tokens, showing that our model achieves 100% accuracy. This test confirms our model’s perfect long-context retrieval capability.
@@
GPT-4-Turbo-2024-04-09 33.16 17.62 19.29 77.73 18.00
   w/ RAG N/A N/A 17.69 77.29 N/A
Qwen2-72B-Instruct 34.88 14.84 21.50 81.66 21.50
   w/ RAG N/A N/A 16.48 76.86 N/A
$\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ 34.11 16.10 44.22 64.63 11.50
   w/ RAG N/A N/A 41.00 70.74 N/A
@@
7.2 Long Context Evaluation Beyond 100K Tokens
In this subsection, we evaluate the long context capability beyond 100K tokens on the real-world tasks from InfiniteBench (Zhang et al., 2024a). Table 2 shows that our model (34.11) outperforms many existing state-of-the-art models, such as GPT4-Turbo-2024-04-09 (33.16), GPT4-1106 preview (28.23), Llama-3-70B-Instruct-Gradient-262k (32.57) and Claude 2 (33.96). Additionally, our model is very close to the highest score of 34.88 achieved by Qwen2-72B-Instruct, confirming the competitive long-context capability of our model.
7.3 Medium-Long Context Evaluation within 32K Tokens
In this subsection, we evaluate the medium-long context capability within 32K tokens. Table 3 shows that GPT-4-Turbo-2024-04-09 achieves the highest score of 51.93 among all models. Our model scores 47.37, which is higher than Llama-3-70B-Instruct-Gradient-262k but is lower than Qwen2-72B-Instruct. This difference can be attributed to the extensive 32K pretraining implemented by Qwen2-72B-Instruct, while we used a much smaller continued pretraining corpus. Additionally, we found that all the RAG solutions perform worse than the long context solution, which suggest all these SOTA long context LLMs can really handle 32K tokens within their context window.
@@ 7.3 Medium-Long Context Evaluation within 32K Tokens
$\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ 47.37 15.20 33.77 81.45 37.27 62.69 53.84
   w/ RAG 47.08 14.71 33.25 80.45 39.45 61.04 53.58
@@
Table 3: Evaluation results on the medium-long context benchmarks within 32K tokens. For RAG, we use top-5 retrieved chunks, each with 1200 tokens from E5-mistral retriever (Wang et al., 2023c).
Models Avg. w/o HDial D2D QuAC QReCC CoQA DoQA CFQA SQA TCQA INSCIT Llama2-Chat-70B 44.64 36.87 32.47 49.40 80.41 38.97 46.85 37.62 44.31 34.88 Llama3-Instruct-70B 52.95 37.88 36.96 51.34 76.98 41.24 76.60 69.61 49.72 36.23 Command R+ 51.40 33.51 34.16 49.77 69.71 40.67 71.21 74.07 53.77 35.76 GPT-3.5-Turbo-0613 50.69 34.83 37.17 50.46 79.33 41.11 73.15 60.63 44.30 35.27 GPT-4-0613 54.35 34.16 40.29 52.01 77.42 43.39 81.28 79.21 45.09 36.34 Llama3-ChatQA-1.5-70B 57.14 41.26 38.82 51.40 78.44 50.76 81.88 83.82 55.63 32.31 GPT-4-Turbo-2024-04-09 54.72 35.35 40.10 51.46 77.73 41.60 84.16 79.98 48.32 33.75 Llama-3-70B-Instruct-Gradient-262k 45.20 34.30 24.01 49.60 73.45 25.76 54.70 63.80 46.30 34.89 Qwen2-72B-Instruct 54.06 35.76 38.48 51.21 85.04 33.89 77.52 77.06 51.64 35.90 $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ 54.81 40.76 38.99 47.12 72.44 51.21 78.52 78.15 55.75 30.35
Table 4: Evaluation results on ChatRAG Bench with 9 datasets. Following (Liu et al., 2024), we exclude HDial as it is included in the instruction tuning datasets. The maximum context lengths are 4K tokens.
@@ 7.4 ChatRAG Bench: Short Context Evaluation within 4K Tokens
In this subsection, we evaluate the models on the short context tasks within 4K tokens from ChatRAG Bench (Liu et al., 2024). Our model achieves the average score of 54.81. Even though it is worse than Llama3-ChatQA-1.5-70B, it still outperforms GPT-4-Turbo-2024-04-09 and Qwen2-72B-Instruct. This confirms that extending short context models to long context is not a free lunch. How to effectively extend the context window to even larger scale (e.g., million tokens in Gemini 1.5 Pro (Gemini-Team, 2024)) without any degradation on regular short context tasks is an exciting research direction.
Top-k 5 10 20 30 Long Context
$\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ 47.08 47.13 47.18 47.19 47.37
Table 5: We compare our $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ using RAG vs. direct long context evaluation on benchmarks with maximum 32K tokens inputs. RAG can be slightly worse than direct long context solution even when we increase top-k chunks to 30.
RAG (top-k) Long Context
$\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ 56.36 (5) 54.43
Qwen2-72B-Instruct 52.95 (20) 51.58
@@
7.5 RAG vs. Long Context
In Table 5 and Table 6, we compare RAG vs. long context solutions under different context lengths. For sequence length beyond 100k, we only report the average score of En.QA and En.MC as the RAG setting is not directly applicable for En.Sum and En.Dia. We found that for downstream tasks within 32k sequence length, our long context solution is better than RAG. This means using RAG can save the cost but the accuracy will drop a bit. On the other hand, we found that for context lengths beyond 100K, RAG (using top-5 for our $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$, and top-20 for Qwen2-72B-Instruct) outperforms the full long-context solution. This indicates that even state-of-the-art long-context LLMs may struggle to effectively understand and reason over 128K tokens. In such scenarios, RAG is recommended for better accuracy and lower inference cost, provided it is applicable to the downstream tasks.
@@ 8 Conclusion
We introduce $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$, a long context model that possess GPT-4 Turbo-level capabilities for both understanding up to 128K long contexts and utilizing retrieved contexts for generation. This provides the flexible options for different downstream tasks with specific accuracy and efficiency requirements. We present a detailed and reproducible technical recipe for building and evaluating the model, including the methods, training data, and evaluation benchmarks. In particular, we evaluate ChatQA 2 on RAG and short-context benchmark (ChatRAG) (within 4K tokens), medium-context tasks from SCROLLS and LongBench (within 32K tokens), and long-context tasks from InfiniteBench (beyond 100K tokens). We demonstrate that the $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ can achieve GPT-4-Turbo-2024-0409 level accuracy on these benchmarks.

### v3

@@ 4 Baselines and Evaluation Benchmarks
We compare our models against SOTA long context models: 1) GPT-4-Turbo-2024-04-09 (128K context window) (OpenAI, 2023), 2) Qwen2-72B-Instruct (128K context window) (Alibaba-QWen, 2024), 3) Llama-3-70B-Instruct-Gradient-262k (256K context window) (GradientAI, 2024), 4) Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct (128K context window). Note that, the Llama-3.1 models are released after this work.
Retrieval-augmented generation (RAG)
We apply RAG to our ChatQA 2 models and baseline long-context LLMs when it is applicable to the downstream tasks, such as those with informative prompts or queries. By default, we use the E5-mistral retriever (Wang et al., 2023c), as it supports both short and long contexts. In our experiments, we use a chunk size of 1200 and select the top 5 chunks, as this configuration provides a good trade-off between cost and performance, as demonstrated in the ablation study in Section 6.
@@
To give a comprehensive study of different context lengths, our evaluation benchmarks covers three categories, 1) ultra-long context benchmarks beyond 100K tokens, 2) long context benchmarks within 32K tokens, and 3) short context benchmarks within 4K tokens.
4.1 Ultra-long Context Benchmarks Beyond 100K Tokens
@@
InfiniteBench (Zhang et al., 2024b) is proposed to evaluate the ultra-long context capability of LLMs over 100K sequence length. As we focus on real-world english tasks, we only take the four related tasks from the InfiniteBench, i.e. longbook summarization (En.Sum), longbook qa (En.QA), longbook multiple choice (En.MC), and longbook dialogue (En.Dia). En.Sum is a task that requires models to generate a concise summary of the given novel and is evaluated using the ROUGE-L-Sum metric (Lin, 2004). En.QA is annotated by a pipeline that ensures the questions’ necessitating of long-range dependencies and reasoning, beyond simple short passage retrieval. Aggregation reasoning and filtering reasoning are the two primary reasoning categories. F1 score is used to evaluate the quality of the answer. En.MC is annotated with the same pipeline of En.QA except that four answer choices are provided and exact matching scores are reported. En.Dia leverages movie and drama scripts from a designated online database  [^4] https://imsdb.com with long, multi-role dialogues. Exact matching score is used again to evaluate the prediction accuracy.
4.2 Long Context Benchmarks within 32K Tokens
@@
We use the long context datasets (except NarrativeQA as it is included in our training) from Xu et al. (2024) as our benchmark for long datasets within 32K. There are six datasets in total, where QMSum (QM), Qasper (QASP), QuALITY (QLTY) are token from SCROLLS (Shaham et al., 2022) and HotpotQA (HQA) MuSiQue (MSQ), MultiFieldQA-en (MFQA) are token from LongBench (Bai et al., 2023). Following the official metrics, we report the geometric mean of ROUGE scores (i.e., ROUGE1/2/L) (Lin, 2004) for QM, the exact matching (EM) score for QLTY, and F1 scores for the remaining four datasets QASP, MSQ, HQA and MFQA.
@@ 4.3 Short Context within 4K Tokens
We use ChatRAG Bench (Liu et al., 2024) as our benchmark for short context within 4k. ChatRAG bench consists of 10 datasets and we exclude HDial as it is included in our training. Following the setup of Liu et al. (2024), for Doc2Dial (D2D), QuAC, and QReCC task with long documents, each document is divided into segments of roughly 300 words. The top 5 relevant chunks are then retrieved as context for each user question. For TopiOCQA and INSCIT, top-20 chunks were retrieved to obtain similar context length to the first three datasets. The other four datasets are CoQA, DoQA, ConvFinQA (CFQA), and SQA, which cover a wide range of domains like finance, children’s stories, literature, mid/high school exams, news, Wikipedia and etc. We use F1 score as the metric to evaluate the generations and report the average score without HDial as it is a fair zero-shot comparisons over different models.
@@ 5 Results
(a) Llama3.1-Instruct 8B
(b) Llama3.1-Instruct 70B
(c) Llama3-ChatQA-2-8B
(d) Llama3-ChatQA-2-70B
Figure 1: Needle In A Haystack test for (a) Llama3.1-Instruct 8B, (b) Llama3.1-Instruct-70B, (c) Llama3-ChatQA-2-8B, and (d) Llama3-ChatQA-2-70B, up to 128K context window. We show the result using the same needle: "The best thing to do in San Francisco is eat a sandwich and sit in Dolores Park on a sunny day."
@@ 5.1 Needle In A Haystack
We evaluate our $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ and $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}8B}$ model on the Needle In A Haystack (NIAH) test (Kamradt, 2023). This synthetic task is popular for testing the long-context capability of LLMs, and can be considered as a threshold level evaluation. We use the needle test: “The best thing to do in San Francisco is eat a sandwich and sit in Dolores Park on a sunny day.” We find this is generally a harder pressure test than the “passkey” in a haystack test, which is commonly used by other long-context LLM works. Figure 3 demonstrates the performance of our models up to 128K tokens. It shows that our model achieves 100% accuracy under both test cases. This test confirms our model’s excellent long-context retrieval capability.
In addition, we also include the NIAH test for Llama3.1-Instruct 8B and Llama3.1-Instruct-70B. From Figures 1(a) and 1(b), we find that neither model could pass the NIAH test with the needle: “The best thing to do in San Francisco…”, although they could pass the easier Passkey retrieval test, as shown in Figures 4(a) and 4(b) in Appendix.
@@
[…]
   w/ RAG N/A N/A 15.41 64.63 N/A
Llama3.1-70B-Instruct 39.81 30.94 38.49 75.55 14.29
   w/ RAG N/A N/A 23.44 79.04 N/A
$\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}8B}$ 35.59 17.14 43.51 64.19 17.50
   w/ RAG N/A N/A 41.48 63.76 N/A
$\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ 41.04 16.08 48.22 80.35 19.50
   w/ RAG N/A N/A 40.66 73.80 N/A
@@
5.2 Ultra-long Context Evaluation Beyond 100K Tokens
In this subsection, we evaluate the ultra-long context capability beyond 100K tokens on the real-world tasks from InfiniteBench (Zhang et al., 2024a). Table 2 shows that our model (41.04) outperforms many existing state-of-the-art models, such as GPT4-Turbo-2024-04-09 (33.16), Llama3.1-70B-Instruct (39.81), Qwen2-72B-Instruct (39.77), Llama-3-70B-Instruct-Gradient-262k (32.57) and Claude 2 (33.96), confirming the effectiveness of our recipe and state-of-the-art long-context capability of our model. We find that our model performs extremely well at QA tasks while performs relatively low for summarization. This can be attributed to the lack of summarization data in our SFT recipe. Additionally, we compare $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}8B}$ to Llama3.1-8B-Instruct and our model again shows better results. The RAG setting shows it is worse than the full long context scores, when we employed the default RAG evaluation setting: top-5 retrieved chunks with a chunk size of 1200 tokens, so overall $5\times 1,200=6,000$ are fed into LLMs. In Section 6 and Table 5, we show that RAG scores can be better than full long context solutions if we increase the number of chunks in the prompt.
5.3 Long Context Evaluation within 32K Tokens
In this subsection, we evaluate the long context capability within 32K tokens. Table 3 shows that GPT-4-Turbo-2024-04-09 achieves the highest score of 51.93 among all models. Our model scores 48.15, which is higher than Llama-3-70B-Instruct-Gradient-262k but is a bit lower than Qwen2-72B-Instruct (49.94) and Llama3.1-70B-Instruct (49.92). This difference can be attributed to the extensive 32K pretraining and large SFT data implemented by Qwen2-72B-Instruct and Llama3.1-70B-Instruct, while we used a much smaller continued pretraining corpus and SFT datasets.
Additionally, we found that the a default RAG solution (top-5 with chunk size as 1200) perform worse than the top long context solution (e.g. GPT-4-Turbo-2024-04-09, Qwen2-72B-Instruct, $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$), which suggests that SOTA long context LLMs can handle 32K tokens well within their context window.
@@ 5.3 Long Context Evaluation within 32K Tokens
[…]
   w/ RAG 41.35 14.08 35.05 65.35 28.30 50.76 54.54
Llama3.1-70B-Instruct 49.92 14.85 35.75 87.40 45.65 59.64 56.23
   w/ RAG 47.63 13.59 32.11 85.65 39.53 58.30 56.58
$\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}8B}$ 39.41 11.64 28.85 63.35 27.81 53.81 51.02
   w/ RAG 40.31 13.20 28.85 61.10 29.77 57.81 51.15
$\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ 48.15 15.99 32.84 82.20 39.26 62.24 56.39
   w/ RAG 47.59 16.16 33.27 81.70 38.16 61.13 55.14
Table 3: Evaluation results on the long context benchmarks within 32K tokens. For RAG, we use top-5 retrieved chunks, each with 1200 tokens from E5-mistral retriever (Wang et al., 2023c). Note that here, we do not intend to optimize the RAG scores by increasing the number of total retrieved tokens and we leave the ablation in Section 6.
[…]
Table 4: Evaluation results on ChatRAG Bench with 9 datasets. Following Liu et al. (2024), we exclude HDial as it is included in the instruction tuning datasets. The maximum context lengths are 4K tokens.
@@ 5.4 ChatRAG Bench: Short Context Evaluation within 4K Tokens
In this subsection, we evaluate the models on the short context tasks within 4K tokens from ChatRAG Bench (Liu et al., 2024). The results can be found in Table 4. Our model achieves an average score of 56.30, outperforming most existing SOTA models with 128K context window, including GPT-4-Turbo-2024-04-09, Qwen2-72B-Instruct, and Llama3.1-70B-Instruct. It is also comparable, though slightly behind, Llama3-ChatQA-1.5-70B (57.14), which only supports 4K context length. This suggests that extending short-context models to long-context has trade-offs. Exploring how to effectively extend the context window to even larger scales, e.g., millions of tokens as in Gemini 1.5 Pro (Gemini-Team, 2024), without degrading performance on regular short-context tasks is an exciting research direction.
[…]
RAG (top-k) Long context $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ 64.55 (k=30) 64.29 Qwen2-72B-Instruct 52.95 (k=20) 52.22
@@
5.5 Ablation Study
Retriever Long (32K) En.MC En.QA E5-Mistral 47.59 73.80 40.66 NV-emb-v2 46.43 74.24 41.19
Table 6: Ablation of different retrievers for $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ using the top-5 and chunk size as 1200.
To understand the choice of different retrievers, we also report the performance of using NV-emb-v2 for $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$, the top-1 retriever in the MTEB leaderboard  [^5] https://huggingface.co/spaces/mteb/leaderboard. Table 6 shows (i) there are relatively small performance variations when switching to another high-performing retriever and (ii) there is no clear overall winner, with NV-Embed-v2 showing better performance on ultra-long tasks.
All-in-one-stage Three-stage Short (4K) 48.55 52.50 Long (32K) 40.69 39.41 Ultra-long (>100K) 35.07 35.59
Table 7: Comparison between the three-stage training and all-in-one-stage training.
To study the effect of three stage recipe, we blended all the instruction-tuning data from all three stages and trained a new model based on Llama-3-8B base in a single stage. The results are shown in Table 10: we observed that the three-stage training process outperforms the all-in-one stage training on RAG by a significant margin, while showing marginally better or worse performance on ultra-long and long tasks.
To further improve the performance of our long (32K) setting, we collect another SFT data which contains 1.6M samples covering a wide range topics. Table 8 shows that our $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}8B}$ (new) is comparable to Llama3.1-8B-Instruct on Long (32K) tasks while being significantly better at Short (4K) and Ultra-long (>100K) tasks. Furthermore, it shows that (i) our model outperforms Llama3.1-8B-Instruct on the math benchmark GSM8K (ii) it underperforms on the knowledge-intensive MMLU and the coding benchmark HumanEval. Since it has not been aligned by DPO / PPO process as Llama3.1-8B-Instruct, the MT-Bench score is lower. We could apply the RLHF process to improve its MT-Bench score in future.
$\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}8B}$ (new) Llama3.1-8B-Instruct Short (ChatRAG) 53.79 48.79 Long (32K) 42.05 42.42 Ultra long (>100K) 35.18 33.17 HumanEval 66.46 70.73 MMLU 65.73 67.59 MT-bench 8.09 8.42 GSM8K (0-shot) 87.41 83.70
Table 8: Comparison between Llama3.1-8B-Instruct and a new $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ model trained by 1.6M SFT dataset
6 RAG versus Long Context
In Figure 2 and Table 5, we compare direct long-context and RAG solutions on 32K and beyond 100K benchmarks. For sequences beyond 100K tokens, we report only the average scores for En.QA and En.MC, as the RAG setting is not directly applicable to En.Sum and En.Dia.
Figure 2 compares different chunk sizes for top-k retrieval and the total number of tokens fed into our long context LLM. Comparing total tokens from 3000 to 24000, we found that more tokens consistently yield better results, confirming the strong long-context capability of our model. Also, for downstream tasks with sequence lengths up to 32K, RAG can still marginally outperform the long-context solution if a sufficient number of top-k chunks are used, i.e., at least 12,000 tokens as shown in Figure 2. Therefore, the hyperparameter k can be selected based on the tradeoff between efficiency and accuracy in real-world applications.
For context lengths beyond 100K in Table 5, RAG (using top-30 chunks for our $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$, and top-20 for Qwen2-72B-Instruct) again outperforms the full long-context solution. In such scenarios, RAG is recommended for better accuracy and much lower inference cost, provided it is applicable to the downstream tasks.
@@ 7 Conclusion
We introduce $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$, a long context LLM that possess GPT-4 Turbo-level capabilities for both understanding up to 128K long contexts and utilizing retrieved contexts for generation. This provides the flexible options for different downstream tasks with specific accuracy and efficiency requirements. We present a detailed and reproducible technical recipe for building and evaluating the model, including the methods, training data, and evaluation benchmarks. In particular, we evaluate ChatQA 2 on short-context RAG benchmark (ChatRAG) (within 4K tokens), long context tasks from SCROLLS and LongBench (within 32K tokens), and ultra-long context tasks from InfiniteBench (beyond 100K tokens). Our results demonstrate that the $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ can achieve GPT-4-Turbo-2024-04-09 level accuracy on these benchmarks across different sequence lengths.
Appendix A Appendix
(a) Llama3.1-Instruct 8B (passkey retrieval)
(b) Llama3.1-Instruct 70B (passkey retrieval)
(c) Llama3-ChatQA-2-8B (passkey retrieval)
(d) Llama3-ChatQA-2-70B (passkey retrieval)
Figure 3: Needle in A Haystack test for Llama3.1-Instruct 8B and Llama3.1-Instruct-70B up to 128K context window. We show the Passkey retrieval results here. The needle is set as "The pass key is 385243. Remember it. 385243 is the pass key." with the question asking "What is the pass key?"
To further confirm the superrior retrieval performance, we evaluate our models on the three retrieval based tasks in InfiniteBench in Table 9. Aligned with our superior NIAH scores, all our models get 100% accuracy for number_string and passkey tasks and our $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ get a better score than GPT-4-Turbo on kv_retrieval task.
model kv_retrieval numbe_string passkey
GPT-4-Turbo 89 100 100
$\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ 91 100 100
$\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}8B}$ 72 100 100
Table 9: $\mathtt{Llama3\text{-}ChatQA\text{-}2\text{-}70B}$ demonstrates better performance on three retrieval tasks in InfiniteBench.
Appendix B Appendix
We trained our model on 2B tokens with different separators and the NIAH below shows that using document breaker “<s>” is much better than <EOS> <BOS>.
(a) NIAH test using <s> as the separator)
(b) NIAH test using <EOS> <BOS> as the separator)
Figure 4: NIAH shows that using document breaker “<s>” is much better than <EOS> <BOS>
Appendix C Appendix
The three-stage SFT process involves blending multiple datasets, with each dataset potentially being trained for a different number of epochs, depending on its blending ratio and number of samples. For instance, the statistics for the datasets used in Stage 2 are provided below, which adds complexity to the analysis.
number of samples epochs
drop 29195 0.68
Narrative_QA 39440 0.69
Quoref 10996 0.68
ROPES 10924 0.69
squad1.1 86863 0.32
squad2.0 86092 0.32
newsqa 76560 0.36
convqa 5413 3.73
nvolve_qa 55458 1.20
scale_cqa 18224 2.37
instruct_v2 1411 4.50
hybriddial 19190 2.70
convfinqa 1162 6.70
tatqa 3176 5.45
corner_cases 1720 4.19
Table 10: Comparison between the three-stage training and all-in-one-stage training.